The perimeter of a equilateral triangle with the side length equal to ____ inches is the same as the perimeter of a rectangle with the length of 15 inches and a width of 12 inches
Answer by jim_thompson5910(35256)

Let's find the perimeter of the rectangle.

P = 2*(L + W)

P = 2*(15 + 12)

P = 2*(27)

P = 54

The perimeter of the rectangle is 54 inches. Because the equilateral triangle has the same perimeter as the rectangle, we know that the triangle has a perimeter of 54 inches.

Perimeter of Rectangle = 54 inches

Perimeter of Equilateral Triangle= 54 inches


-------------------------------------------------------

We'll take this perimeter of 54 inches and divide it evenly in 3 equal pieces. Remember that all equilateral triangles have 3 equal sides.

54/3 = 18

So this triangle has each side length equal to 18 inches.

Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim